OpenStudy (anonymous):

1,3,7,15,31 what two sequence next this numbers

OpenStudy (anonymous):

twice previous +1

OpenStudy (anonymous):

3-1=2 7-3=4 15-7=8 31-15=16 x-31=32 ..Therefore x= 63 Similarly continue